Waterway Construction Company has entered into a contract beginning January 1, 2020, to build a parking complex. It has been estimated that the complex will cost $598,000 and will take 3 years to construct. The complex will be billed to the purchasing company at $897,000. The following data pertain to the construction period.

2020 2021 2022

Costs to date $275,080 $412,620 $607,000

Estimated costs to complete 322,920 185,380 –0–

Progress billings to date 272,000 545,000 897,000

Cash collected to date 242,000 495,000 897,000

(a) Using the percentage-of-completion method, compute the estimated gross profit that would be recognized during each year of the construction period. (If answer is 0, please enter 0. Do not leave any fields blank.)

Gross profit recognized in 2020  

Gross profit recognized in 2021

Gross profit recognized in 2022  

(b) Using the completed-contract method, compute the estimated gross profit that would be recognized during each year of the construction period. (If answer is 0, please enter 0. Do not leave any fields blank.)

Gross profit recognized in 2020

Gross profit recognized in 2021  

Gross profit recognized in 2022
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Waterway Construction Company
Ans to a
Particulars 2020 2021 2022 Note
Contract Price             897,000             897,000             897,000 A
Cost Incurred during the year             275,080             137,540             194,380 B
Estimated cost to complete             322,920             185,380                        -   C
Cost Incurred to date             275,080             412,620             607,000 D
(Sum total of the cost incurred during the year)
Estimated total cost             598,000             598,000             607,000 E=C+D
Percentage complete 46.00% 69.00% 100.00% F=D/E
(Cost Incurred to date/Estimated total cost)
Revenue to be recognized             412,620             206,310             278,070 G
(Contract Price*Percentage complete)- revenue previously recognized
Year 2020: 46% completed. Revenue recognized = 46% x $ 897,000 = $ 412,620.
Year 2021: 69% completed. Revenue recognized = 69% x $ 897,000 – $ 412,620 (previously recognized) = $ 206,310.
Year 2022: 100% completed. Revenue recognized = 100% x $ 897,000 – $ 412,620- $ 206,310 (previously recognized) = $ 278,070.
Gross Profit/(Loss)             137,540               68,770               83,690 H=G-B
Ans to b
Particulars 2020 2021 2022 Note
Contract Price             897,000             897,000             897,000 A
Cost Incurred during the year             275,080             137,540             194,380 B
Cost Incurred to date             275,080             412,620             607,000 D
Revenue to be recognized                        -                          -               897,000 G
Gross Profit/(Loss)                        -                          -               290,000 H=G-D